# FEDONet : Fourier-Embedded DeepONet for Spectrally Accurate Operator Learning

### v2

@@ 1 Introduction
Partial differential equations underpin models ranging from turbulent‐flow dynamics and heat diffusion to electromagnetics and biological transport. Classical numerical solvers such as finite‐difference, finite‐element and spectral methods provide high‐fidelity solutions but incur a prohibitive computational burden when faced with high‐dimensional, multiscale or stiff systems [1, 2, 3]. Efforts to construct efficient surrogate models have a long history, spanning the universal approximation theorems for neural networks[4, 5, 6, 7], reduced‐order techniques based on proper‐orthogonal decomposition [8, 9], kernel Gaussian process surrogates [10], and mesh‐free radial‐basis schemes [11, 12]. However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long standing challenge.
The recent emergence of operator learning addresses this long-standing challenge by training neural networks that act directly on infinite-dimensional function spaces, enabling generalization across discretizations and grid resolutions. Deep Operator Networks [13] pioneered the branch–trunk architecture for learning nonlinear operators, while subsequent Neural Operator formulations [14, 15, 16] extended these ideas with rigorous operator–approximation guarantees and scalable architectures. Building on these foundations, Fourier Neural Operators [17] employ global spectral convolutions to capture nonlocal correlations efficiently, achieving remarkable speed-ups and accuracy gains in parametric PDE learning. Subsequent developments have further advanced this paradigm through multiwavelet representations [18, 19], convolutional kernels [20], multigrid tensorizations [21, 22], and geometry-aware spectral deformations [23], collectively enhancing scalability, resolution fidelity, and geometric adaptability.
@@
Despite the recent operator-learning architectures, the trunk network of conventional DeepONets remains a shallow multilayer perceptron. This design often struggles to represent the oscillatory and multi-scale structures that characterize nonlinear PDE solutions, resulting in reduced sample efficiency and reconstruction accuracy. While recent alternatives such as graph-based [26] or wavelet-enhanced U-Net trunks [28] offer improved expressivity, they introduce significant architectural complexity and domain-specific preprocessing requirements. In contrast, random Fourier features have proven remarkably effective at encoding high-frequency information in standard deep networks [48], yet their potential within operator-learning frameworks remains largely unexplored, where the trunk is typically implemented as a simple fully connected MLP. To address this gap, we propose the Fourier-Embedded Deep Operator Network (FEDONet), which augments the DeepONet trunk by replacing raw spatial coordinates with a compact random Fourier-feature embedding [48]. This embedding implicitly introduces a spectrum of sinusoidal basis functions, enriching the trunk’s representational capacity with negligible additional parameters. FEDONet retains the original branch–trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
We formally introduce the Fourier embedding of trunk coordinates and analyze its theoretical implications through the lens of the operator neural tangent kernel, demonstrating that the embedding strictly enlarges the approximation class of standard DeepONets. A comprehensive benchmark is carried spanning seven canonical PDE families - steady Poisson, viscous Burgers, Allen–Cahn, Kuramoto–Sivashinsky, Eikonal, Lorenz–63, and Lorenz–96. By embedding a principled spectral bias into the trunk network, FEDONet bridges the performance gap to frequency-domain operators, while preserving the architectural simplicity, flexibility, and locality benefits of the original DeepONet formulation.
Section 2 presents the formulation of the Fourier-embedded Deep Operator Network (FEDONet), detailing the embedding construction, its theoretical connection to the operator neural tangent kernel, and its implications for approximation capacity. Section 3 describes the benchmark setup and implementation details across the seven PDE families considered. Section 4 reports quantitative and spectral analyses of FEDONet performance, including convergence behavior, and comparisons with baseline operator models. Finally, Section 5 summarizes the main findings and outlines directions for future research in spectral operator learning and physics-constrained network design.
@@ 2.1 Fourier Embeddings as Spectral Preconditioners
$\phi(\zeta)=\left[\sin(2\pi B\zeta),\cos(2\pi B\zeta)\right],\quad B_{ij}\sim\mathcal{N}(0,\sigma^{2}),$ (4)
where $B\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ - enriching its spectral content.
@@ 3 Results
Figure 2: Comparison of predicted and reference solution fields for the two-dimensional Poisson equation.
Figure 2 shows a representative test case from the Poisson dataset. The top row depicts the input forcing, reference solution, prediction from the baseline DeepONet, and the corresponding pointwise error, while the bottom row presents results from the proposed Fourier-Embedded DeepONet (FEDONet). FEDONet reproduces the reference field with sharper gradients and more accurate spatial localization of extrema, whereas the baseline model exhibits noticeable blurring and displacement of high-magnitude regions. The residual field of the vanilla DeepONet fluctuates within $\pm 1.5\times 10^{-3}$ and displays fine-scale artifacts, while FEDONet reduces the range to $\pm 3\times 10^{-4}$ and yields a smoother, spatially coherent error distribution, corresponding to an approximate fivefold improvement in reconstruction accuracy.
Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ (Fig. 3). Both models reproduce the reference spectrum at large scales ($k\lesssim 5$), but differences arise in the intermediate and dissipative ranges. The baseline DeepONet consistently overestimates energy for $k\gtrsim 8$, resulting in a spurious buildup of high-wavenumber content that mirrors the oscillatory residuals seen in physical space. In contrast, FEDONet closely follows the reference spectrum across the entire range, deviating by less than a factor of two near the Nyquist limit and maintaining a physically consistent energy decay.
@@
Taken together, the spatial and spectral analyses demonstrate that the Fourier embedding enhances the trunk network’s ability to capture high-frequency features and improves overall stability. FEDONet achieves lower pointwise errors while eliminating the artificial spectral plateau that characterizes conventional fully connected trunks.
@@ 3.2 Burgers’ Equation
The Burgers’ equation is a well-known 1D viscous equation. It has been studied extensively for DeepONets in prior literature. The equation is given by:
@@
Figure 4 shows the relative $\ell_{2}$ errors over 128 unseen test samples. FEDONet consistently outperforms DeepONet, with a mean error reduction from 4.01% to 2.63%, and a narrower spread in error distribution indicating increased reliability.
Figure 5 presents the best-performing test sample, showing the ground truth solution alongside predictions from both models. While DeepONet captures the coarse solution structure, it struggles with steep gradients and exhibits visible smoothing near shock regions. In contrast, FEDONet produces significantly sharper and more accurate reconstructions, faithfully capturing shock fronts and fine-scale dynamics.
Figure 4: Burgers Relative $\ell_{2}$ error across 128 unseen samples for DeepONet and FEDONet.
The energy spectra of the predicted and ground truth fields, shown in Figure 6, further emphasize this advantage. FEDONet accurately captures the spectral decay across a wide range of wavenumbers, aligning with the physical expectations of dissipative transport phenomena. DeepONet, however, underestimates energy at higher modes, reflecting its inability to resolve fine features and spectral richness. This demonstrates that the Fourier embeddings effectively bridge the model’s spectral bias, enabling it to better represent high-frequency dynamics essential to Burgers-type flows.
Figure 5: Best-performing test sample for Burgers’ Equation: Ground truth vs. DeepONet vs. FEDONet prediction.
These results underscore the importance of aligning model inductive bias with the physical spectrum of the target PDE. The enhanced performance of FEDONet in both physical and spectral domains highlights its potential as a more robust and generalizable operator learner for convection-dominated systems.
Figure 6: Temporal Evolution of Pointwise Error and Energy Spectrum for Burgers’ Equation
@@
Figure 7: Training loss convergence for the 1D Burgers’ Equation - Comparison of Vanilla and Fourier Embedded DeepONets trained with varying batch sizes.
3.3 Lorenz–63 System
The Lorenz–63 system is a widely studied chaotic dynamical system that serves as a benchmark for evaluating the long-term stability and predictive accuracy of surrogate models. It is governed by the following system of coupled nonlinear ordinary differential equations:
@@ 3.3 Lorenz–63 System
Figure 8: Reconstructed Lorenz–63 attractor for a randomnly sampled initial condition
@@
Figure 9: Phase-plane projection of the xxzz states. The vanilla DeepONet trajectory diverges from the expected orbit, showing geometric drift and incorrect lobe transitions. FEDONet reconstructs the spiral arms and inter-lobe dynamics with higher accuracy, preserving the overall topology.
@@
Figure 10: Pointwise prediction error in $x(t)$, $y(t)$, and $z(t)$ for a representative sample. The vanilla DeepONet exhibits compounding errors, especially after $t>1.5$, reflecting the accumulation of approximation drift. FEDONet achieves significantly lower errors and better error containment beyond lobe transitions, highlighting improved stability under chaotic evolution.
@@
Figure 11: Cumulative relative error over time. DeepONet exhibits rapid exponential error growth beyond $t>1.5$, characteristic of chaotic divergence. FEDONet accumulates error more gradually, indicating improved long-term stability and mitigation of runaway prediction drift.
@@
Figure 12: Spectral energy density of predicted $x(t)$ trajectories compared with the ground truth. DeepONet underestimates energy in mid-to-high frequency bands, resulting in oversmoothing and loss of chaotic detail. FEDONet preserves the spectral tail and matches the broadband structure more accurately, contributing to improved phase and amplitude resolution.
@@ 3.4 Eikonal Equation
A total of 1250 distinct airfoil shapes are synthesized by randomly sampling parameters $m\in[0.01,0.09]$, $p\in[0.1,0.7]$, and $t\in[0.1,0.4]$. This dataset captures a wide range of geometrical variations and curvature profiles, providing a rigorous benchmark for learning PDE solution operators from geometric input.
Figure 13: Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry for Eikonal Equation. Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
@@
3.5 Lorenz-96 Chaotic System
To assess model performance on nonlinear dynamical systems, we consider the Lorenz-96 (L96) model, a prototypical chaotic system widely used for testing data-driven surrogates in climate modeling and nonlinear forecasting. The governing equations are
$\frac{dx_{i}}{dt}=(x_{i+1}-x_{i-2})x_{i-1}-x_{i}+F,\quad i=1,2,\dots,N,$ (26)
subject to periodic boundary conditions $x_{i-N}=x_{i}=x_{i+N}$. Here, $x_{i}$ denotes the state of the $i$-th variable, $N$ is the system dimension, and $F$ is a constant forcing parameter that determines the level of nonlinearity and chaos. For $F>8$, the system exhibits strong chaotic behavior; in this study, we set $F=4.0$ to obtain moderately nonlinear yet stable dynamics.
The system is integrated using a classical fourth-order Runge–Kutta (RK4) scheme with a fixed time step $\Delta t=0.01$:
$\mathbf{x}_{n+1}=\mathbf{x}_{n}+\frac{\Delta t}{6}(k_{1}+2k_{2}+2k_{3}+k_{4}),$ (27)
where $k_{1},k_{2},k_{3},k_{4}$ are the intermediate RK stages computed from the Lorenz–96 derivatives.
A dataset of 10,00010{,}000 independent trajectories is generated for training and evaluation. Each trajectory is initialized as
$\mathbf{x}_{0}=F\cdot\mathbf{1}+\epsilon\,\mathcal{N}(0,I),\quad\epsilon=10^{-3},$
with $N=40$ and integrated for a total physical time of 15 s. The initial 10 s are discarded as a transient, and the remaining 5 s (501 time steps) are retained to capture quasi-stationary behavior. The resulting dataset has dimensions (10,000,501,40)(10{,}000,501,40), representing 10,000 spatio-temporal trajectories of 40-dimensional state vectors.
Figure 14: Lorenz-96 - Complete spatio-temporal field reconstruction comparison between DeepONet and FEDONet.
Figure 14 shows representative test trajectories from the L96 dataset. Each panel depicts the initial condition, reference evolution, predictions from the baseline DeepONet and the proposed Fourier-Embedded DeepONet (FEDONet), and their respective absolute-error fields. Both models reproduce the characteristic diagonal stripe patterns associated with wave-like phase propagation, maintaining amplitude and wavelength stability over the full 500-step forecast horizon. No visible drift or blow-up is observed in either solution.
Quantitatively, the residual fields indicate comparable performance. In the first trajectory, the vanilla DeepONet attains a root-mean-square (RMS) error on the order of $\mathcal{O}(10^{-1})$, while FEDONet slightly reduces this magnitude; in the second trajectory, the ordering reverses locally, but the overall difference remains small. On average, FEDONet achieves marginally lower global $\ell_{2}$ errors, though the improvement is less pronounced than in the Poisson or Burgers benchmarks.
This outcome reflects the inherent structure of the L96 dynamics: the solution manifold is dominated by a slowly varying, low-dimensional spatio-temporal mode, which a shallow fully connected trunk can already approximate effectively. Consequently, introducing random Fourier features provides limited additional benefit, consistent with the quantitative results in Table 1, where both models achieve sub-25% relative error.
@@ 3.6 Allen–Cahn Equation
We evaluate the performance of the proposed Fourier-Embedded DeepONet (FEDONet) in modeling the one-dimensional Allen–Cahn equation nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems. The governing partial differential equation is given by:
@@
Figure 15: Best-performing test case for the Allen–Cahn equation. Relative $\ell_{2}$ error: $4.85\%$ for DeepONet, $2.27\%$ for FEDONet.
Figure 15 shows qualitative predictions for the best-performing test case. FEDONet sharply resolves the interface zones and phase transitions, while DeepONet produces smoother approximations with notable blurring near discontinuities. The spectral inductive bias introduced by randomized Fourier features enables FEDONet to retain high-frequency details crucial to phase-field dynamics. The pointwise absolute error plot in Figure 16 emphasizes the localized improvements of FEDONet in capturing steep interfaces. DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
Figure 16: Allen-Cahn - Energy Spectrum and Pointwise absolute error across space and time for a representative sample.
Figure 16 compares the spatial energy spectra of predicted and true solutions. FEDONet accurately tracks the ground truth spectrum across all wavenumbers, preserving the high-frequency tail associated with sharp phase boundaries. In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen–Cahn system. These results collectively demonstrate that the incorporation of spectral priors via Fourier embeddings in FEDONet enhances both pointwise accuracy and spectral fidelity. The improved resolution of phase interfaces and better retention of high-frequency information position FEDONet as a robust surrogate model for nonlinear, multiscale PDEs such as the Allen–Cahn equation.

### v3

@@ Abstract
Deep Operator Networks (DeepONets) have recently emerged as powerful data-driven frameworks for learning nonlinear operators, particularly suited for approximating solutions to partial differential equations. Despite their promising capabilities, the standard implementation of DeepONets, which typically employs fully connected linear layers in the trunk network, can encounter limitations in capturing complex spatial structures inherent to various PDEs. To address this limitation, we introduce Fourier-Embedded trunk networks within the DeepONet architecture, leveraging random fourier feature mappings to enrich spatial representation capabilities. Our proposed Fourier-Embedded DeepONet, FEDONet demonstrates superior performance compared to the traditional DeepONet across a comprehensive suite of PDE-driven datasets, including the two-dimensional Poisson, Burgers’, Lorenz-63, Eikonal, Allen-Cahn, and the Kuramoto-Sivashinsky equation. FEDONet delivers consistently superior reconstruction accuracy across all benchmark PDEs, with particularly large relative $L^{2}$ error reductions observed in chaotic and stiff systems. This study highlights the effectiveness of Fourier embeddings in enhancing neural operator learning, offering a robust and broadly applicable methodology for PDE surrogate modeling.
keywords:
Neural Operator Learning , Fourier Feature Embeddings , Deep Operator Networks , Partial Differential Equations
journal: Journal of Computational Physics
@@
organization=Mechanical and Aerospace Engineering, addressline=1512 Middle Drive, University of Tennessee, city=Knoxville, postcode=37996-2210, state=TN, country=USA
@@ 1 Introduction
Partial differential equations underpin models ranging from turbulent‐flow dynamics and heat diffusion to electromagnetics and biological transport. Classical numerical solvers such as finite‐difference, finite‐element and spectral methods provide high‐fidelity solutions but incur a prohibitive computational burden when faced with high‐dimensional, multiscale or stiff systems [1, 2, 3]. Efforts to construct efficient surrogate models have a long history, spanning the universal approximation theorems for neural networks[4, 5, 6, 7], reduced‐order techniques based on proper‐orthogonal decomposition [8, 9], kernel Gaussian process surrogates [10], and mesh‐free radial‐basis schemes [11, 12]. Early data‐driven approaches also sought to learn governing equations directly from trajectory data using deep neural networks [13] However, extending these approaches to learn function‐to‐function mappings capable of generalizing across discretizations, boundary conditions, and input configurations has remained a long standing challenge.
The recent emergence of operator learning addresses this long-standing challenge by training neural networks that act directly on infinite-dimensional function spaces, enabling generalization across discretizations and grid resolutions. Deep Operator Networks [14] pioneered the branch-trunk architecture for learning nonlinear operators, while subsequent Neural Operator formulations [15, 16, 17] extended these ideas with rigorous operator-approximation guarantees and scalable architectures. Building on these foundations, Fourier Neural Operators [18] employ global spectral convolutions to capture nonlocal correlations efficiently, achieving remarkable speed-ups and accuracy gains in parametric PDE learning. Related ideas have been explored in the context of learning PDE dynamics in modal or Fourier space [19], where the evolution operator is approximated directly on spectral coefficients rather than physical fields, highlighting the efficacy of frequency-space representations for capturing multiscale dynamics. Subsequent developments have further advanced this paradigm through multiwavelet representations [20, 21], convolutional kernels [22], multigrid tensorizations [23, 24], and geometry-aware spectral deformations [25], collectively enhancing scalability, resolution fidelity, and geometric adaptability.
@@
Despite the recent operator-learning architectures, the trunk network of conventional DeepONets remains a shallow multilayer perceptron. This design often struggles to represent the oscillatory and multi-scale structures that characterize nonlinear PDE solutions, resulting in reduced sample efficiency and reconstruction accuracy. While recent alternatives such as graph-based [28] or wavelet-enhanced U-Net trunks [30] offer improved expressivity, they introduce significant architectural complexity and domain-specific preprocessing requirements. In contrast, random Fourier features have proven remarkably effective at encoding high-frequency information in standard deep networks [50], yet their potential within operator-learning frameworks remains largely unexplored, where the trunk is typically implemented as a simple fully connected MLP. To address this gap, we propose the Fourier-Embedded Deep Operator Network (FEDONet), which augments the DeepONet trunk by replacing raw spatial coordinates with a compact random Fourier-feature embedding [50]. This embedding implicitly introduces a spectrum of sinusoidal basis functions, enriching the trunk’s representational capacity with negligible additional parameters. FEDONet retains the original branch-trunk separation, integrates seamlessly with existing DeepONet implementations, and incurs no runtime overhead beyond a single matrix multiplication.
We formally introduce the Fourier embedding of trunk coordinates and analyze its theoretical implications through the lens of the operator neural tangent kernel, demonstrating that the embedding strictly enlarges the approximation class of standard DeepONets. A comprehensive benchmark is carried spanning seven canonical PDE families - steady Poisson, viscous Burgers, Allen-Cahn, Kuramoto-Sivashinsky, Eikonal, and Lorenz-63. By embedding a principled spectral bias into the trunk network, FEDONet bridges the performance gap to frequency-domain operators, while preserving the architectural simplicity, flexibility, and locality benefits of the original DeepONet formulation. Section  2 presents the formulation of the Fourier-Embedded Deep Operator Network (FEDONet), detailing the embedding construction, its theoretical connection to the operator neural tangent kernel, and its implications for approximation capacity. Section  3 describes the benchmark setup and implementation details across the seven PDE families considered. Sections  4 and  5 summarize the main findings and outlines directions for future research in spectral operator learning and physics-constrained network design.
@@ 2.1 Fourier Embeddings as Spectral Preconditioners
$\phi(\zeta)=\left[\sin(2\pi Z\zeta),\cos(2\pi Z\zeta)\right],\quad Z_{ij}\sim\mathcal{N}(0,\sigma^{2}),$ (4)
where $Z\in\mathbb{R}^{M\times d}$ is a Gaussian frequency matrix. Instead of feeding raw coordinates to the trunk, we input $\phi(\zeta)$ - enriching its spectral content.
@@ 3 Results
Figure 2: Solving a 2D Poisson equation: Comparison of predicted and reference solution fields of DeepONet and FEDONet
Figure 2 shows a representative test case from the Poisson dataset. The top row depicts the input forcing field, exact solution, prediction from the baseline DeepONet, and the corresponding pointwise error, while the bottom row presents results from the proposed Fourier-Embedded DeepONet (FEDONet). Both models were evaluated on $1000$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $1.41\%$ for the DeepONet and $1.09\%$ for the FEDONet. FEDONet reproduces the reference field with sharper gradients and more accurate spatial localization of extrema. The residual field of the vanilla DeepONet displays fine-scale artifacts, while FEDONet yields a smoother, spatially coherent error distribution, corresponding to an approximate fivefold improvement in reconstruction accuracy.
@@
Spectral accuracy was assessed using the angle-integrated energy spectrum $E_{k}$ shown in Fig. 3. Both models reproduce the reference spectrum at large scales ($k\lesssim 50$), but differences arise in the intermediate and dissipative ranges. The baseline DeepONet overestimates energy for $k\gtrsim 90$, resulting in a spurious buildup of high-wavenumber content that mirrors the oscillatory residuals seen in physical space. In contrast, FEDONet closely follows the reference spectrum across the entire range. Taken together, the spatial and spectral analyses demonstrate that the Fourier embedding enhances the trunk network’s ability to capture high-frequency features and improves overall stability. FEDONet achieves lower pointwise errors while eliminating the artificial spectral plateau that characterizes conventional fully connected trunks.
@@ 3.2 Burgers’ Equation
The Burgers’ equation is a well-known 1D viscous equation. It has been studied extensively for DeepONets in prior literature [53]. The equation is given by:
@@
Both models were evaluated on $250$ unseen test samples, yielding average relative $L^{2}$ errors of $6.43\%$ for the DeepONet and $4.86\%$ for the FEDONet. Figure 4 presents prediction for a test sample, showing the ground truth solution alongside predictions from both models. While DeepONet captures the coarse solution structure, it struggles with steep gradients and exhibits visible smoothing near shock regions. In contrast, FEDONet produces significantly sharper and more accurate reconstructions, faithfully capturing shock fronts and fine-scale dynamics.
Figure 4: Solving a parametric Burgers’ equation: Comparison of predicted spatio-temporal field for a random initial condition. Relative $L^{2}$ error: $4.09\%$ for DeepONet, $1.49\%$ for FEDONet.
The energy spectra of the predicted and ground truth fields, shown in Figure 5, further emphasize this advantage. FEDONet accurately captures the spectral decay across a wide range of wavenumbers, aligning with the physical expectations of dissipative transport phenomena. DeepONet, however, underestimates energy at higher modes, reflecting its inability to resolve fine features and spectral richness. This demonstrates that the Fourier embeddings effectively bridge the model’s spectral bias, enabling it to better represent high-frequency dynamics essential to Burgers-type flows. The enhanced performance of FEDONet in both physical and spectral domains highlights its potential as a more robust and generalizable operator learner for convection-dominated systems.
Figure 5: Energy spectrum and pointwise absolute error across space and time for a representative sample for Burgers’ Equation.
@@
Figure 6: Training loss convergence for the 1D Burgers’ Equation - Comparison of DeepONet and FEDONet trained with varying batch sizes.
3.3 Lorenz-63 System
The Lorenz-63 system is a widely studied chaotic dynamical system that serves as a benchmark for evaluating the long-term stability and predictive accuracy of surrogate models. It is governed by the following system of coupled nonlinear ordinary differential equations:
@@ 3.3 Lorenz-63 System
Figure 7: Reconstructed Lorenz-63 attractor for a randomnly sampled initial condition for trained DeepONet and FEDONet
@@
Figure 8: Phase-plane projection of the xxzz states.
@@
Figure 9: Pointwise prediction error in $x(t)$, $y(t)$, and $z(t)$ for a representative sample.
@@
Figure 10: Cumulative relative error over time.
@@
Figure 11: Spectral energy density of predicted $x(t)$ trajectories compared with the ground truth.
@@ 3.4 Eikonal Equation
A total of 1250 distinct airfoil shapes are synthesized by randomly sampling parameters $m\in[0.01,0.09]$, $p\in[0.1,0.7]$, and $t\in[0.1,0.4]$. Of these, 1000 samples are used for training and 250 for testing. This dataset captures a wide range of geometrical variations and curvature profiles, providing a rigorous benchmark for learning PDE solution operators from geometric input. On this benchmark, the baseline DeepONet attains an average relative $L^{2}$ error of 2.15%, while the proposed FEDONet achieves an improved 1.12% error, demonstrating substantially higher reconstruction accuracy.
Figure 12: Solving a parametric Eikonal equation (airfoils): Comparison of predicted signed distance functions (SDFs) for a representative airfoil geometry. Relative $L^{2}$ error: $2.728\times 10^{-2}$ for DeepONet, $0.912\times 10^{-2}$ for FEDONet.
@@ 3.5 Allen–Cahn Equation
We further evaluate the performance of the proposed FEDONet in modeling the one-dimensional Allen–Cahn nonlinear reaction-diffusion system widely used to describe phase separation processes in multi-component systems. The governing partial differential equation is given by:
@@
Figure 13: Solving a parametric Allen Cahn equation: Comparison of predicted spatio-temporal field for a random initial condition. Relative $L^{2}$ error: $4.85\%$ for DeepONet, $2.27\%$ for FEDONet.
Both models were evaluated on $128$ randomly selected unseen test samples, yielding average relative $L^{2}$ errors of $10.61\%$ for the DeepONet and $5.50\%$ for the FEDONet. Figure 13 shows qualitative predictions for a representative test case. FEDONet sharply resolves the interface zones and phase transitions, while DeepONet produces smoother approximations with notable blurring near discontinuities. The pointwise absolute error plot in Figure 14 emphasizes the localized improvements of FEDONet in capturing steep interfaces. DeepONet accumulates substantial errors in regions with high spatial gradients, whereas FEDONet maintains predictive accuracy throughout the domain, suggesting better handling of dissipative nonlinearities.
Figure 14: Energy spectrum and pointwise absolute error across space and time for a representative sample for Allen-Cahn Equation.
Figure 14 compares the spatial energy spectra of predicted and true solutions. FEDONet accurately tracks the ground truth spectrum across all wavenumbers, preserving the high-frequency tail associated with sharp phase boundaries. In contrast, DeepONet suffers from spectral attenuation at $k>20$, revealing an inability to fully capture dissipative and oscillatory dynamics inherent in the Allen-Cahn system.